OpenStudy (anonymous):

leah is 6 years older than sue john is 5 years older than leah their totals = 41 how old is sue

OpenStudy (anonymous):

ok. So let's say that leah's age is the variable x. then john must be x + 5 right? since john is 5 years old then leah. and you are given the total sum of theirages which is 41. so you can rwite john + leah = 41 which is equivalent to x+5 + x = 41 ... 2x + 5 = 41 and this equation can be solved now. So 2x = 41 -5 = 36 therefore x = 18 so leah's age is 18. hence john's age is 18 + 5 = 23. and if you double check 23 + 18 it adds upto 41! :)

OpenStudy (anonymous):

ok one moment i am going in the military so all this work has to be done without a calculator and me and math do not mix well

OpenStudy (anonymous):

how did you get 2x

OpenStudy (anonymous):

ok no problem. think about problems similar to these in this manner: you don't know the age of leah or john right? HOWEVER, we are given the relation between the two. and that relation is that john is 5 years older. so now you can say since we dont know the age of leah, let's just say its some random number - so we say let it be equal to X. since we dont know what X is. then we say ok so we know age of leah is X. therefore using the rlation of john, john's age must be x +5 right? since he's 5 years older. now you are given the sum of their ages. so the age of leah + age of john must be 41. so we can say that x (leahs age) + 5+x (johns age) = 41. so x + x + 5 = 41. Now whenever you add variables, we just sum them up. so say if we knew that x = 2. then 2 + 2 = 4 right? similarly 2*2 is also 4 right? so we can ssay using the same logic that since we have x + x = some number then x + x = 2 * x. makes sense? and then you can solve the equation 2x + 5 = 41 like i did before.

OpenStudy (anonymous):

hang on one min, is sue, john and leahs ages all equal 41?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

uthere

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

ok so if sue's age is x, then leah's age must be x+6 since she is 6 years older . then it says john is 5 years older than leah. so compared to leah whos 6 years older then sue, then john must be 11 years older than sue. so john is x+11. so from here we know that their total sum is 41. so we can say that: x + (x+6) + (x+11) = 41. soo x+ x+ x+ 17 = 41 Hence 3x = 24. and x = 8. Therefore sue's age must be 8. then leah's age is 8+6 = 14. and john is 8+11 = 19.